# Peeling Context from Cause for Multimodal Molecular Property Prediction

### v1

@@ C.2 FreeSolv Dataset
Figure 8: Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s hydrophilicity the most.
@@
Figure 9: Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s hydrophilicity the most.
At the optimal depth (L2), CLaP assigns the highest weights to the carbonyl oxygen and the exocyclic amine group, with additional contributions from the ring nitrogens. The 3D branch aggregates the localized dipoles into a polar spine, while the 2D branch captures the carbonyl and amino substituents from connectivity. This matches physical chemistry: the carbonyl oxygen and multiple NH2/NH groups provide hydrogen-bond acceptors and donors, substantially raising aqueous solubility relative to a purely carbocyclic core. Under-peeling diffuses weights across the ring scaffold (shortcuts), while over-peeling smears attribution and reintroduces residual context. Heteroaryl amides of this type are commonly found in nucleobases and drug-like scaffolds, where their balance of hydrogen-bonding capacity and planarity helps tune solubility in chemical biology.
C.3 Lipo Dataset
@@
Figure 10: Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s hydrophilicity the most.
At the optimal depth (L3), CLaP assigns the highest weights to the hydroxyl substituent and nearby heteroaryl nitrogens. The 3D branch aggregates the O–H dipole into a localized polar hotspot, while the 2D branch highlights the hydroxyl group and its attachment point to the heteroaromatic ring. This depth-wise focus is exactly the intended “peel”: batch-coupled context is removed, and label-relevant structure (polar donors/acceptors distributed across a largely aromatic scaffold) is retained. The outcome matches physical chemistry: the hydroxyl provides a strong hydrogen-bond donor and acceptor site, partially counterbalancing the extended aromatic surface. As a result, the molecule is predicted to be more hydrophilic than a fully unsubstituted aromatic analog.
Figure 11: Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s hydrophilicity the most.
At the optimal depth (L3), CLaP assigns the highest weight to the amide carbonyl oxygen and the nearby amide nitrogen. The 3D branch captures the strong local dipole between the C=O and NH, while the 2D branch highlights the amide linkage from connectivity. This depth-wise focus is exactly the intended “peel”: batch-coupled context is removed, and label-relevant structure (a polar amide motif appended to a hydrophobic ring) is retained. CLaP captures a general modification in chemistry: the amide group generally introduces both a hydrogen-bond donor and acceptor, increasing polarity and reducing lipophilicity relative to an unsubstituted carbocyclic scaffold. As a result, the molecule is predicted to be more hydrophilic (lower $\log D_{7.4}$).
C.4 CycPeptMPDB Dataset
@@
Figure 12: Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s permeability the most.
At the optimal depth (L5), CLaP highlights hydrophobic residues such as phenylalanine and aliphatic side chains as the dominant contributors. The 3D branch aggregates bulky aromatic surfaces into hydrophobic patches, while the HELM representation captures residue-level context. This depth-wise focus is exactly the intended “peel”: batch-coupled noise is removed, and label-relevant features (hydrophobic residues that enhance membrane passage) are retained. The outcome is consistent with physical chemistry and with what is known for peptide therapeutics: aromatic and aliphatic substitutions increase lipophilicity and improve passive cell permeability. More broadly, such analyses underscore the drug-design logic of cyclic and linear peptide scaffolds: by balancing polarity (for solubility) with nonpolar residues (for permeability), peptide-based molecules can be optimized as potential drugs. Peptide drugs already play a major role in therapy, ranging from hormones and enzyme inhibitors to next-generation modulators of protein–protein interactions.
Figure 13: Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s permeability the most.
At the optimal depth (L5), CLaP again highlights hydrophobic residues such as aromatics and aliphatic side chains as the main drivers of permeability. These groups form contiguous lipophilic patches, favoring bilayer partitioning and consistent with design principles for permeable peptides. In contrast, heteroatom-rich sites such as oxygen incorporated cycles contribute less, reflecting their polarity and the desolvation penalty that reduces passive diffusion. Thus, the model recovers a chemically intuitive balance: hydrophobic groups increase permeability, while oxygen and general incorporation of polar atoms in the side chains reduce permeability. CLaP is able to distinguish both drivers of permeability, as well as identify motifs that harm permeability, suggesting molecular edits that chemists can make to optimize the peptide.
@@ Appendix D Implementation Details
D.1 Proposed Framework Implementation Details
For CycPeptMPDB, we encode 2D molecular graphs and HELM strings with GAT and 3D geometries with EGNN; each backbone has 3 layers, producing 128-D embeddings that are linearly projected to a shared 256-D space. The target correlation schedule increases linearly from $0.5$ to $0.8$ across L=5L{=}5 causal blocks. Training runs for 100 epochs with batch size 16 on $8\times$ RTX-6000 GPUs with early stopping. Loss weights are $\lambda_{\mathrm{caus}}{=}1.0$, $\lambda_{\mathrm{unif}}{=}1.0$, and $\lambda_{\mathrm{mono}}{=}0.5$; cross-modal consistency is disabled ($\lambda_{\mathrm{cons}}{=}0$), and the hinge margin is set to 0. For ESOL, FreeSolv, and Lipo (non-peptide datasets), we use two modalities (2D graphs and 3D geometries); unless noted otherwise, embeddings are projected to 256-D and the correlation schedule is $0.4{\rightarrow}0.7$. ESOL uses width 128 with 2 encoder layers and L=2L{=}2 causal blocks; FreeSolv uses width 64 with 2 layers and L=2L{=}2; Lipo uses width 64 with 3 layers and L=3L{=}3. The hinge margin is 0 for all settings.
D.2 Baselines Implementation Details
@@
Appendix E The Use of Large Language Models (LLMs)
We used an LLM solely for light editorial assistance (grammar, wording, and minor style). The LLM did not contribute ideas, methods, models, experiments, analyses, or decisions, and was not used to generate or label data or code. All scientific content is by the authors.

### v2

@@ C.2. FreeSolv Dataset
Figure 9. Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s hydrophilicity the most.
For Figure 10, at the optimal depth (L2), CLaP assigns the highest weights to the carbonyl oxygen and the exocyclic amine group, with additional contributions from the ring nitrogens. The 3D branch aggregates the localized dipoles into a polar spine, while the 2D branch captures the carbonyl and amino substituents from connectivity. This matches physical chemistry: the carbonyl oxygen and multiple NH2/NH groups provide hydrogen-bond acceptors and donors, substantially raising aqueous solubility relative to a purely carbocyclic core. Under-peeling diffuses weights across the ring scaffold (shortcuts), while over-peeling smears attribution and reintroduces residual context. Heteroaryl amides of this type are commonly found in nucleobases and drug-like scaffolds, where their balance of hydrogen-bonding capacity and planarity helps tune solubility in chemical biology.
Figure 10. Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s hydrophilicity the most.
C.3. Lipo Dataset
@@
For Figure 11, at the optimal depth (L3), CLaP assigns the highest weights to the hydroxyl substituent and nearby heteroaryl nitrogens. The 3D branch aggregates the O–H dipole into a localized polar hotspot, while the 2D branch highlights the hydroxyl group and its attachment point to the heteroaromatic ring. This depth-wise focus is exactly the intended “peel”: batch-coupled context is removed, and label-relevant structure (polar donors/acceptors distributed across a largely aromatic scaffold) is retained. The outcome matches physical chemistry: the hydroxyl provides a strong hydrogen-bond donor and acceptor site, partially counterbalancing the extended aromatic surface. As a result, the molecule is predicted to be more hydrophilic than a fully unsubstituted aromatic analog.
Figure 11. Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s hydrophilicity the most.
For Figure 12, at the optimal depth (L3), CLaP assigns the highest weight to the amide carbonyl oxygen and the nearby amide nitrogen. The 3D branch captures the strong local dipole between the C=O and NH, while the 2D branch highlights the amide linkage from connectivity. This depth-wise focus is exactly the intended “peel”: batch-coupled context is removed, and label-relevant structure (a polar amide motif appended to a hydrophobic ring) is retained. CLaP captures a general modification in chemistry: the amide group generally introduces both a hydrogen-bond donor and acceptor, increasing polarity and reducing lipophilicity relative to an unsubstituted carbocyclic scaffold. As a result, the molecule is predicted to be more hydrophilic (lower $\log D_{7.4}$).
Figure 12. Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s hydrophilicity the most.
C.4. CycPeptMPDB Dataset
@@
For Figure 13, at the optimal depth (L5), CLaP highlights hydrophobic residues such as phenylalanine and aliphatic side chains as the dominant contributors. The 3D branch aggregates bulky aromatic surfaces into hydrophobic patches, while the HELM representation captures residue-level context. This depth-wise focus is exactly the intended “peel”: batch-coupled noise is removed, and label-relevant features (hydrophobic residues that enhance membrane passage) are retained. The outcome is consistent with physical chemistry and with what is known for peptide therapeutics: aromatic and aliphatic substitutions increase lipophilicity and improve passive cell permeability. More broadly, such analyses underscore the drug-design logic of cyclic and linear peptide scaffolds: by balancing polarity (for solubility) with nonpolar residues (for permeability), peptide-based molecules can be optimized as potential drugs. Peptide drugs already play a major role in therapy, ranging from hormones and enzyme inhibitors to next-generation modulators of protein–protein interactions.
Figure 13. Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s permeability the most.
For Figure  14, at the optimal depth (L5), CLaP again highlights hydrophobic residues such as aromatics and aliphatic side chains as the main drivers of permeability. These groups form contiguous lipophilic patches, favoring bilayer partitioning and consistent with design principles for permeable peptides. In contrast, heteroatom-rich sites such as oxygen incorporated cycles contribute less, reflecting their polarity and the desolvation penalty that reduces passive diffusion. Thus, the model recovers a chemically intuitive balance: hydrophobic groups increase permeability, while oxygen and general incorporation of polar atoms in the side chains reduce permeability. CLaP is able to distinguish both drivers of permeability, as well as identify motifs that harm permeability, suggesting molecular edits that chemists can make to optimize the peptide.
Figure 14. Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s permeability the most.
@@ Appendix D Implementation Details
D.1. Proposed Framework Implementation Details
Our framework instantiates modality-specific encoders for 2D molecular graphs, sequence representations, and 3D geometries. Specifically, we encode 2D molecular graphs and HELM strings with GAT and 3D geometries with EGNN(3D structures apply ETKDG-up to 5 initial conformers, and energy-minimize the first embedded conformer using UFF); each backbone has 3 layers, producing 128-D embeddings that are linearly projected to a shared 256-D space. The target correlation schedule increases linearly from $0.5$ to $0.8$ across L=5L{=}5 causal blocks. Training runs for 100 epochs with batch size 16 on $8\times$ RTX-6000 GPUs with early stopping. Loss weights are $\lambda_{\mathrm{caus}}{=}1.0$, $\lambda_{\mathrm{unif}}{=}1.0$, and $\lambda_{\mathrm{mono}}{=}0.5$; cross-modal consistency is disabled ($\lambda_{\mathrm{cons}}{=}0$), and the hinge margin is set to 0. Our implementation is publicly available at an anonymous GitHub repository (https://anonymous.4open.science/r/CLaP-FE64).
D.2. Baselines Implementation Details